# Gradient Existence and Energy Finiteness of Local Minimizers in the Wasserstein $L^{\infty}$ Topology for Binary-Star Systems

### v1

@@ 5.3 Minimizers’ existence and nonexistence
$U\left({\rho+t\sigma}\right)\geq{\int_{\{{\rho>N_{1}}\}}{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}}\geq\frac{1}{L+1}\left({1-\frac{t}{\epsilon(\sigma)}}\right){\int_{\{{\rho>N_{1}}\}}{A(\rho)}}=\infty$
@@
Notice that, in Proposition 5.23, we do not require the topology being inherited from a topological vector space. In fact, when defining a weak local minimizer (see Appendix A.1), we do not utilize any specific properties of any topology. (Although we call it “weak” because, under the concept of a topological vector space, a local minimizer can imply a weak local minimizer. See, for example, Lemma 5.14.)
We can also recall Remark 5.10, which states that, in the topology induced by $W^{\infty}$ distance, a local minimizer is not necessarily a weak local minimizer. Therefore, it is interesting to compare Proposition 5.23 and Lemma 5.7.
Appendix A Appendix
Here we introduce some preliminary knowledge which are used in the paper.
A.1 Calculus of Variations in Vector Spaces
@@
Let $X$ be a real vector space, $U$ a subset of $X$, and $E$ an extended real-valued function defined on $U$. Given $u\in U$, let $P(u)=\{v\in X\mid\exists\epsilon(v)>0,\text{ such that }u+tv\in U\text{ for }t\in[0,\epsilon(v))\}$. This set is a cone in $X$. If $P$ is any cone in $X$, let $X(P)$ be the linear subspace of $X$ generated by $P$. We say that $E$ is $P$-differentiable at $u$ if $P\subset P(u)$ and there is a linear functional $\mu$ on $X(P)$ such that $\forall v\in P$
$\lim\limits_{t\rightarrow 0^{+}}t^{-1}\{E(u+tv)-E(u)\}=\mu(v),$
The linearity of $\mu$ is needed in our analysis; its precise domain is not essential here. We denote the “derivative” $\mu$ by $E_{u}^{\prime}$—with $P$ understood from context.
@@
In fact, the latter corresponds to the definition of weak local minimizer in [5, Section 2]. However, this limited-based definition does not fully align with our intuitive understanding of a “minimizer”. For instance, consider $E(u)=-u^{2}$; according to this limited-based definition, $u=0$ would qualify as a weak local minimizer. Yet, this is not a reasonable conclusion. Our definition, on the other hand, helps exclude such case.
@@
Generally speaking, if $X$ is a topological vector space, that is, $u+tv$ is continuous w.r.t. $t$, then one can prove that a local minimizer is a weak local minimizer. For reference, see Lemma 5.14. What’s more, the “weak” nature of a weak local minimizer, compared to a local minimizer, is also reflected in the following:
To check $u$ is weak local minimizer, we just need to show $E(u)$ is a minimum over $E(u+t_{n}v)$ for some $t_{n}$} in any arbitrary direction $v\in P(u)$, rather than for all small $t$ in the direction $v$. Moreover, we do not need to specify the topology or find a neighborhood of $\rho$.
@@
A.2 Properties of Sobolev Spaces
We recall some properties of Sobolev spaces as well as of $L^{p}$ spaces. Some of them will be used frequently in this thesis.
Proposition A.6 (Interpolation Inequality [9, Section 4.2]).
If $f\in L^{p}\cap L^{q}$ with $1\leq p\leq q\leq\infty$, then $f\in L^{r}$ for all $r,p\leq r\leq q$, and the following interpolation inequality holds:
$\|f\|_{r}\leq\|f\|_{p}^{\alpha}\|f\|_{q}^{1-\alpha},\text{ where }\frac{1}{r}=\frac{\alpha}{p}+\frac{1-\alpha}{q},0\leq\alpha\leq 1$
Proposition A.7 (Hardy-Littlewood-Sobolev Inequality [8, Theorem 1.7]).
Let $1<p,r<\infty$ and $0<\alpha<n$ be such that $\frac{1}{p}+\frac{\alpha}{n}=\frac{1}{r}+1$. $\exists C_{p,\alpha,n}>0$, such that
$\left\||\cdot|^{-\alpha}*f\right\|_{L^{r}\left(\mathbb{R}^{n}\right)}\leq C_{p,\alpha,n}\|f\|_{L^{p}\left(\mathbb{R}^{n}\right)}$ (33)
Remark A.8.
Notice that $|x|^{-\alpha}$ is not in $L^{\frac{n}{\alpha}}(\mathbb{R}^{3})$, hence Young’s Inequality [9, Theorem 4.33 and Exercise 4.30] is not applicable. But Hardy-Littlewood-Sobolev Inequality Proposition A.7 can also fail at the case $p=1,r=\infty$, or $\alpha=n$.
Notice that Hardy-Littlewood-Sobolev Inequality Proposition A.7 can fail when $r=\infty$. To estimate the bound in $L^{\infty}$, we introduce the following proposition:
Proposition A.9 (Bound of Potential [5, Proposition 5]).
@@ A.2 Properties of Sobolev Spaces
The last inequality comes from Young’s Inequality [9, Theorem 4.33 and Exercise 4.30] with $\frac{1}{p-\epsilon}+\frac{1}{(p-\epsilon)^{\prime}}=1$ and $\frac{1}{1+\epsilon}+$ $\frac{1}{(1+\epsilon)^{\prime}}=1$. Since $p>\frac{3}{2}$, we can choose a small $\epsilon$ such that those norms are finite. And then apply the Interpolation Inequality Proposition A.6 for $\|\rho\|_{L^{p-\epsilon}}$ amd $\|\rho\|_{L^{1+\epsilon}}$ to obtain (34). Moreover, we can check $\forall x\in\mathbb{R}^{3},V_{\rho}(x)$ is well defined and finite. (The idea is similar as the following proof of continuity.) Let $\tau_{h}f(x)=f(x+h)$, then
$\left|V_{\rho}(x+h)-V_{\rho}(x)\right|=$ $\left|\int\frac{\rho(x+h-y)-\rho(x-y)}{|y|}d^{3}y\right|$
$\leq$ $\left|\int\left(\tau_{h}\rho(x-y)-\rho(x-y)\right)b_{1}(y)d^{3}y\right|$
$+\left|\int\left(\tau_{h}\rho(x-y)-\rho(x-y)\right)b_{2}(y)d^{3}y\right|$
@@
The last inequality comes from the Hölder’s inequality. Notice $p>\frac{3}{2}$ thus its Hölder conjugate $p^{\prime}<3$ thus $\left\|b_{1}\right\|_{L^{p^{\prime}}}<\infty$. $\left\|\tau_{h}\rho-\rho\right\|_{L^{p}}$ and $\left\|\tau_{h}\rho-\rho\right\|_{L^{1}}$ go to 0 when $h\rightarrow 0$. Therefore, $V_{\rho}$ is continuous and $\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is finite. Thus $V_{\rho}$ is bounded, i.e. $\exists C>0,\forall x\in\mathbb{R}^{3},|V_{\rho}(X)|<C$. ∎
@@
Lemma A.10 (Continuous Differentiable Representative of Potential).
If $V\in W^{1,\infty}\left(\mathbb{R}^{3}\right)$, and its weak derivative and itself are continuous. Then $V$ has a representative which is continuously differentiable and weak derivative coincides with the classical one a.e. in $\mathbb{R}^{3}$.
@@
$V\in W^{1,\infty}\left(\mathbb{R}^{3}\right)$. Therefore, $\forall R>0,V$ has a representative which is Lipschitz in $B_{R}(0)$ [15, Subsection 5.8.2], and then for almost every $x\in\mathbb{R}^{3},V$ is differentiable and the weak derivative coincides with the classical one by Rademacher’s theorem [15, Subsection 5.8.3]. To replace “almost everywhere” by “everywhere”, we modify the discussion about Sobolev space in one dimension in [9, Section 8.2]. Without loss of generality, we assume $j=1$, and consider $\widetilde{f}\left(x_{1},x_{2},x_{3}\right)=\int_{0}^{x_{1}}\frac{\partial V}{\partial x_{1}}\left(y,x_{2},x_{3}\right)dy$, where $\frac{\partial V}{\partial x_{1}}$ is continuous by assumption, one can show its weak and classical derivative with respect to $x_{1}$ are $\frac{\partial V}{\partial x_{1}}$ both in $\mathbb{R}^{3}$ and in $\mathbb{R}$. Therefore, $g:=\widetilde{f}-V$ is continuous and has 0 as its weak partial derivative. Consider $g_{n}=\rho_{n}*g$, where $\left\{\rho_{n}\right\}$ is a sequence of mollifiers, the classical derivative $\frac{\partial g_{n}}{\partial x_{1}}=\frac{\partial\rho_{n}}{\partial x_{1}}*g$ [9, Proposition 4.20], while by the definition of weak derivative we have $\frac{\partial\rho_{n}}{\partial x_{1}}*g=\int\frac{\partial\rho_{n}}{\partial x_{1}}(x-y)g(y)dy=\int\rho_{n}(x-y)\frac{\partial g}{\partial x_{1}}(y)dy=0$. Therefore, for all point in $\mathbb{R}^{3},g_{n}\left(x_{1},x_{2},x_{3}\right)=c_{n}\left(x_{2},x_{3}\right)$. Because $g_{n}\rightarrow g$ uniformly on compact sets of $\mathbb{R}^{3}$ [9, Proposition 4.21], in particular converges everywhere, $g\left(x_{1},x_{2},x_{3}\right)=$ $c\left(x_{2},x_{3}\right)$ for some function $c$. Therefore, $\widetilde{f}\left(x_{1},x_{2},x_{3}\right)=c\left(x_{2},x_{3}\right)+V_{\rho}\left(x_{1},x_{2},x_{3}\right)$ everywhere in $\mathbb{R}^{3}$. Then we take the classical derivative with respect to $x_{1}$ again to see $V_{\rho}$’s classical derivative is $\frac{\partial V_{\rho}}{\partial x_{1}}$, which is also continuous. ∎
It can be shown in the distribution sense we have $\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int\frac{y_{j}\rho(x-y)}{|y|^{3}}d^{3}y$. Therefore, combine the arguments and statements as above we have:
Proposition A.11 (Differentiability of Potential [5, Proposition 7]).
@@
Similar to the arguments in proof of Proposition A.9, we know $\frac{\partial V_{\rho}}{\partial x_{j}}\in L^{\infty}\left(\mathbb{R}^{3}\right)\cap C\left(\mathbb{R}^{3}\right)$, and together with Proposition A.9 we have $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)\cap C\left(\mathbb{R}^{3}\right)$. Therefore, by Lemma A.10 we get the results. ∎
Remark A.12.
One can also prove Proposition A.11 by direct computation.
@@ Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez (master’s thesis advisor) and Dimitri Cobb (second advisor). Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.

### v2

@@ 5.3 Minimizers’ existence and nonexistence
$U\left({\rho+t\sigma}\right)\geq{\int_{\{{\rho>N_{1}}\}}{A\left({\left({1-\frac{t}{\epsilon(\sigma)}}\right)\rho}\right)}\,dx}\geq\frac{1}{L+1}\left({1-\frac{t}{\epsilon(\sigma)}}\right){\int_{\{{\rho>N_{1}}\}}{A(\rho)}\,dx}=\infty$
@@
Notice that, in Proposition 5.21, we do not require the topology to be inherited from a topological vector space. In fact, when defining a weak local minimizer (see Appendix A), we do not utilize any specific properties of any topology.
Appendix
Here we introduce some preliminary knowledge that is used in the paper.
Appendix A Calculus of Variations in Vector Spaces
@@
Let $X$ be a real vector space, $U$ a subset of $X$, and $E$ an extended real-valued function defined on $U$. Given $u\in U$, let $P(u)=\{v\in X\mid\exists\epsilon(v)>0,\text{ such that }u+tv\in U\text{ for }t\in[0,\epsilon(v))\}$. This set is a cone in $X$. If $P$ is any cone in $X$, let $X(P)$ be the linear subspace of $X$ generated by $P$. We say that $E$ is $P$-differentiable at $u$ if $P\subset P(u)$ and there is a linear functional $l$ on $X(P)$ such that $\forall v\in P$
$\lim\limits_{t\rightarrow 0^{+}}t^{-1}\{E(u+tv)-E(u)\}=l(v),$
The linearity of $l$ is needed in our analysis; its precise domain is not essential here. We denote the “derivative” $l$ by $E_{u}^{\prime}$—with $P$ understood from context.
@@
In fact, the latter corresponds to the definition of weak local minimizer in [5, Section 2]. However, this limited-based definition does not fully align with our intuitive understanding of a “minimizer”. For instance, consider $E(u)=-u^{2}$; according to this limit-based definition, $u=0$ would qualify as a weak local minimizer. Yet, this is not a reasonable conclusion. Our definition, on the other hand, helps exclude such case.
@@
Generally speaking, if $X$ is a topological vector space, that is, $u+tv$ is continuous w.r.t. $t$, then one can prove that a local minimizer is a weak local minimizer. See, for example, Lemma 5.13. This is also why we use the term “weak”.
To check $u$ is a weak local minimizer, we just need to show $E(u)$ is a minimum over $E(u+t_{n}v)$ for some $t_{n}$} in any arbitrary direction $v\in P(u)$, rather than for all small $t$ in the direction $v$. Moreover, we do not need to specify the topology or find a neighborhood of $\rho$. See also Remark 5.22.
@@
Appendix B Properties of Sobolev Spaces
We review some properties of Sobolev spaces and $L^{p}$ spaces, complemented by proofs for certain propositions. Some of them will be used frequently in this paper.
First we notice that Hardy-Littlewood-Sobolev Inequality [8, Theorem 1.7] can fail when we consider the $L^{\infty}$ norm of the convolution. To estimate the bound in $L^{\infty}$, we introduce the following proposition:
Proposition B.1 (Bound of Potential [5, Proposition 5]).
@@ Appendix B Properties of Sobolev Spaces
The last inequality comes from Young’s Inequality [9, Theorem 4.33 and Exercise 4.30] with $\frac{1}{p-\epsilon}+\frac{1}{(p-\epsilon)^{\prime}}=1$ and $\frac{1}{1+\epsilon}+$ $\frac{1}{(1+\epsilon)^{\prime}}=1$. Since $p>\frac{3}{2}$, we can choose a small $\epsilon$ such that those norms are finite. And then apply the Interpolation Inequality [9, Section 4.2] for $\|\rho\|_{L^{p-\epsilon}}$ amd $\|\rho\|_{L^{1+\epsilon}}$ to obtain (34). Moreover, we can check $\forall x\in\mathbb{R}^{3},V_{\rho}(x)$ is well defined and finite. (The idea is similar to that of the following proof of continuity.) Let $\tau_{h}f(x)=f(x+h)$, then
$\left|V_{\rho}(x+h)-V_{\rho}(x)\right|=$ $\left|\int_{\mathbb{R}^{3}}\frac{\rho(x+h-y)-\rho(x-y)}{|y|}\,dy\right|$
$\leq$ $\left|\int_{\mathbb{R}^{3}}\left(\tau_{h}\rho(x-y)-\rho(x-y)\right)b_{1}(y)\,dy\right|$
$+\left|\int_{\mathbb{R}^{3}}\left(\tau_{h}\rho(x-y)-\rho(x-y)\right)b_{2}(y)\,dy\right|$
@@
The last inequality comes from Hölder’s inequality. Notice $p>\frac{3}{2}$ thus its Hölder conjugate $p^{\prime}<3$ thus $\left\|b_{1}\right\|_{L^{p^{\prime}}}<\infty$. $\left\|\tau_{h}\rho-\rho\right\|_{L^{p}}$ and $\left\|\tau_{h}\rho-\rho\right\|_{L^{1}}$ go to 0 when $h\rightarrow 0$. Therefore, $V_{\rho}$ is continuous and $\left\|V_{\rho}\right\|_{L^{\infty}\left(\mathbb{R}^{3}\right)}$ is finite. Thus $V_{\rho}$ is bounded, i.e. $\exists C>0,\forall x\in\mathbb{R}^{3},|V_{\rho}(X)|<C$. ∎
@@
Lemma B.2 (Continuous Differentiable Representative of Potential).
If $V\in W^{1,\infty}\left(\mathbb{R}^{3}\right)$, and its weak derivative and itself are continuous. Then $V$ has a representative which is continuously differentiable and the weak derivative coincides with the classical one a.e. in $\mathbb{R}^{3}$.
@@
$V\in W^{1,\infty}\left(\mathbb{R}^{3}\right)$. Therefore, $\forall R>0,V$ has a representative which is Lipschitz in $B_{R}(0)$ [15, Subsection 5.8.2], and then for almost every $x\in\mathbb{R}^{3},V$ is differentiable and the weak derivative coincides with the classical one by Rademacher’s theorem [15, Subsection 5.8.3]. To replace “almost everywhere” by “everywhere”, we modify the discussion about Sobolev space in one dimension in [9, Section 8.2]. Without loss of generality, we assume $j=1$, and consider $\widetilde{f}\left(\widetilde{x_{1}},x_{2},x_{3}\right)=\int_{0}^{x_{1}}\frac{\partial V}{\partial x_{1}}\left(y,x_{2},x_{3}\right)\,d\widetilde{x_{1}}$, where $\frac{\partial V}{\partial x_{1}}$ is continuous by assumption, one can show its weak derivative and classical derivative with respect to $x_{1}$ are $\frac{\partial V}{\partial x_{1}}$ both in $\mathbb{R}^{3}$ and in $\mathbb{R}$. Therefore, $g:=\widetilde{f}-V$ is continuous and has 0 as its weak partial derivative. Consider $g_{n}=\rho_{n}*g$, where $\left\{\rho_{n}\right\}$ is a sequence of mollifiers, the classical derivative $\frac{\partial g_{n}}{\partial x_{1}}=\frac{\partial\rho_{n}}{\partial x_{1}}*g$ [9, Proposition 4.20], while by the definition of weak derivative we have $\frac{\partial\rho_{n}}{\partial x_{1}}*g(x)=\int_{\mathbb{R}^{3}}\frac{\partial\rho_{n}}{\partial x_{1}}(x-y)g(y)dy=\int_{\mathbb{R}^{3}}\rho_{n}(x-y)\frac{\partial g}{\partial x_{1}}(y)dy=0$. Therefore, for all point in $\mathbb{R}^{3},g_{n}\left(x_{1},x_{2},x_{3}\right)=c_{n}\left(x_{2},x_{3}\right)$. Because $g_{n}\rightarrow g$ uniformly on compact sets of $\mathbb{R}^{3}$ [9, Proposition 4.21], in particular converges everywhere, $g\left(x_{1},x_{2},x_{3}\right)=$ $c\left(x_{2},x_{3}\right)$ for some function $c$. Therefore, $\widetilde{f}\left(x_{1},x_{2},x_{3}\right)=c\left(x_{2},x_{3}\right)+V_{\rho}\left(x_{1},x_{2},x_{3}\right)$ everywhere in $\mathbb{R}^{3}$. Then we take the classical derivative with respect to $x_{1}$ again to see $V_{\rho}$’s classical derivative is $\frac{\partial V_{\rho}}{\partial x_{1}}$, which is also continuous. ∎
Let $h(x)=-\int_{\mathbb{R}^{3}}\frac{y_{j}\rho(x-y)}{|y|^{3}}dy$. Similar to the arguments in proof of Proposition B.1, we know $h$ is in $L^{q}(\mathbb{R}^{3})$ for $\rho\in L^{1}\left(\mathbb{R}^{3}\right)\cap L^{p}(\mathbb{R}^{3})$ under appropriate indices $p$ and $q$. Consequently, through the use of test functions and Fubini’s theorem, it follows that $h$ is the weak derivative of $V_{\rho}$, i.e., the following result holds in the sense of distributions:
$\frac{\partial V_{\rho}}{\partial x_{j}}(x)=-\int_{\mathbb{R}^{3}}\frac{y_{j}\rho(x-y)}{|y|^{3}}\,dy.$
Therefore, combining the arguments and statements as above, we have:
Proposition B.3 (Differentiability of Potential [5, Proposition 7]).
@@
Similar to the arguments in proof of Proposition B.1, we know $\frac{\partial V_{\rho}}{\partial x_{j}}\in L^{\infty}\left(\mathbb{R}^{3}\right)\cap C\left(\mathbb{R}^{3}\right)$, and together with Proposition B.1 we have $V_{\rho}\in$ $W^{1,\infty}\left(\mathbb{R}^{3}\right)\cap C\left(\mathbb{R}^{3}\right)$. Therefore, by Lemma B.2 we get the results. ∎
Remark B.4.
One can also prove Proposition B.3 by direct computation.
@@ Acknowledgments
The author is partially supported by the National Science Foundation grant DMS-2308208. This work was primarily carried out during the author’s Master’s studies at the University of Bonn. The author thanks Juan Velázquez and Dimitri Cobb for their continued advice and support since the author’s time in Bonn, as well as Christof Sparber and Mimi Dai for their comments and support during the author’s Ph.D. studies at the University of Illinois Chicago. Thanks also to Lorenzo Pompili, Shao Liu, Xiaopeng Cheng, Bernhard Kepka, Daniel Sánchez Simón del Pino for discussions, and to Théophile Dolmaire and other instructors. The author is grateful to his parents.